### v2

# A recursive approach for the determination of the nice sections of width three which have a 4-crown stack as retract Version 8-3

Abstract
The characterization of the finite minimal automorphic posets of width three is still an open problem. Niederle has shown that this task can be reduced to the characterization of the nice sections of width three having a non-trivial tower of nice sections as retract. In our article, we develop a recursive approach for the determination of those of them which have a 4-crown stack as retract. We apply the approach on a sub-class $\mathfrak{N}_{2}$ of nice sections of width three and determine all posets in $\mathfrak{N}_{2}$ with height up to six having a 4-crown stack as retract. For each integer $n\geq 2$, the class $\mathfrak{N}_{2}$ contains $2^{n-2}$ different isomorphism types of posets of height $n$.
Mathematics Subject Classification:
Primary: 06A07. Secondary: 06A06.
Key words: fixed point, fixed point property, retract, retraction, minimal automorphic, width three.
1 Introduction
We call a finite poset $P$ automorphic if it has a fixed point free automorphism, and we call it minimal automorphic if additionally every proper retract of it has the fixed point property. Minimal automorphic posets have been present in the literature since the late eighties; an overview is provided by [10]. In particular, they have found interest as forbidden retracts because a poset does not have the fixed point property iff it has a minimal automorphic poset as retract [9, Th. 4.8].
However, the description or characterization of minimal automorphic posets is still an open field. (Unless otherwise stated, “poset” always means “finite poset” in this article.) Since the pioneering work of Rival [7, 8], we know that a poset of height one does not have the fixed point property iff it contains a crown, because a crown of minimal length will be a retract of it. Brualdi and Dias Da Silva [1] generalized this result by showing that a poset does not have the fixed point property iff it has a “generalized crown” as retract. With respect to posets of small width, Fofanova and Rutkowski [4] showed that a poset of width two does not have the fixed point property iff it has a 4-crown stack as retract. (The result extends even to chain-complete infinite posets.) Niederle [5] characterized in 1989 posets of width three not having the fixed point property by having a non-trivial retract being a “tower of nice sections”. Later on, in 2008, he extended his result to posets of width four [6].
Unfortunately, it is quite difficult to break down Niederle’s abstract characterization to concrete posets or classes of posets. As far as we know, the only real success in this field is a result of Farley [3] from 1997, who showed that the ranked nice sections of width three are exactly the 6-crown stacks, and that a 6-crown stack is minimal automorphic iff its height $h$ is not a multiple of three. Otherwise, it has a 4-crown stack of height $\frac{2}{3}h$ as retract. Schröder [10, Lemma 2.6] generalized one direction of this result by showing that a $2n$-crown stack of height $nm$ has a 4-crown stack of height $2m$ as retract. Moreover, in the case of $n\geq 3$, a $2n$-crown stack of height $h$ is not minimal automorphic if $h$ and $n$ have a non-trivial common factor [10, Prop. 2.7].
Due to the results of Niederle [5], the study of the minimal automorphic posets of width three can be restricted to the investigation of the nice sections of width three. In our article, we develop and apply a recursive approach for their determination. After the preparatory Section 2, we deal with the classes $\mathfrak{N}$ of nice sections and $\mathfrak{C}$ of crowned sections in Section 3; $\mathfrak{N}$ is a proper subclass of $\mathfrak{C}$. We see that a poset $P\in\mathfrak{C}$ has a 4-crown stack as retract iff it has such a retract containing a split level, and we prove several results about the existence of such levels. We see in Lemma 3.11 that a split level of $R$ corresponds in $P$ to a certain decomposition by a retractive split, and based on this observation we develop at the end of Section 3 a recursive approach for the determination of the nice sections having a 4-crown stack as retract.
In order to test the approach, we introduce the class $\mathfrak{N}_{2}$ of nice sections with horizon two in Section 4. For each integer $n\geq 2$, this class contains $2^{n-2}$ isomorphism types of posets of height $n$; the isomorphism types are described in Theorem 4.2. We characterize in Theorem 4.5 the retractive splits of posets $P\in\mathfrak{N}_{2}$ yielding a 4-crown stack as retract. Using this result, we determine in Section 5 by means of the recursive approach all posets $P\in\mathfrak{N}_{2}$ of height up to six which have a 4-crown stack as retract.
The characterization of the minimal automorphic posets of width three is regarded as a difficult task [9, p. 98], in parts even as “intractable” [3, p. 126]. Also the sparse sequence of publications indicates these difficulties. Nevertheless, it gives some confidence that the nice section of width three used by Schröder [9, p. 99] for the illustration of the problems and considered as “nasty” by Farley [3, p. 136] is handled without a problem by the machinery developed in this article. It is the poset coded as 1011 in Section 5 and shown in Figure 9.
2 Preparation
In this section, we introduce our notation and recapitulate definitions of structures being in the focus of our investigation. For all other terms of order theory, the reader is referred to standard textbooks as [9]. For integers $i\leq j$, we write $[i,j]$ for the interval $\{i,\ldots,j\}$.
Let $P=(X,\leq)$ be a finite poset. (We deal with finite posets only in this paper.) For $Y\subseteq X$, the induced sub-poset $P|_{Y}$ of $P$ is $\left(Y,{\leq}\cap(Y\times Y)\right)$. To simplify notation, we identify a subset $Y\subseteq X$ with the poset $P|_{Y}$ induced by it.
For $y\in X$, we define the down-set and up-set induced by $y$ as
${\downarrow\,}y$ $:=\left\{x\in X\>\left|\>x\leq y\right.\right\},\quad{\downarrow^{{}^{\!\!\!\!\!\!\circ}}}y:=({\downarrow\,}y)\setminus\{y\},$
${\uparrow\,}y$ $:=\left\{x\in X\>\left|\>y\leq x\right.\right\},\quad{\uparrow_{{}_{{}_{\!\!\!\!\!\!\circ}}}}y:=({\uparrow\,}y)\setminus\{y\}.$
Given two points $x<y$ in $P$, the point $x$ is called a lower cover of $y$ and $y$ an upper cover of $x$ if $x\leq z\leq y$ implies $z\in\{x,y\}$ for all $z\in P$; this relation is denoted by $x\lessdot y$. A point in $P$ is called irreducible if it has a single lower cover or a single upper cover.
Let $A,B\subseteq P$. We write $A<B$ if $a<b$ for all $a\in A$, $b\in B$. If $A=\{a\}$ is a singleton, we simply write $a<B$ and $B<a$. The notation for the relation $\leq$ is analogous.
The length of a chain is its cardinality minus 1, and the height of a poset is the maximal length of a chain contained in it. For a poset $P$, we denote its height by $h_{P}$. The width of a poset is the largest size of an antichain contained in it.
For a poset $P$, the level sets $P(k)$, $k\in[0,h_{P}]$, are recursively defined by
$P(0)$ $:=\min P,$
$P(k+1)$ $:=\min\left(P\setminus\cup_{i=0}^{k}P(i)\right)\quad\>\textrm{for all }\>0\leq k<h_{P}.$
Additionally, we write for all $k,\ell\in[0,h_{P}]$ with $k\leq\ell$,
$P(k,\ell):=P(k)\cup P(\ell)\quad\>\textrm{and}\>\quad P(k\rightarrow\ell):=\cup_{i=k}^{\ell}P(i).$
For level indices $k_{1}\leq k_{2}$ of $P$, we call the sub-poset $P(k_{1}\rightarrow k_{2})$ a segment of $P$, a lower segment if $k_{1}=0$, and an upper segment if $k_{2}=h_{P}$.
For $n\in\mathbb{N}\setminus\{1\}$, a $2n$-crown is a poset with $2n$ points $x_{i,j}$, $i\in[0,n-1]$, $j\in\{0,1\}$, in which $x_{0,0}<x_{0,1}>x_{1,0}<\cdots>x_{n-1,0}<x_{n-1,1}>x_{0,0}$ are the only comparabilities between different points. We call a poset $P$ of height at least 1 a $2n$-crown stack if $P(k,k+1)$ is a $2n$-crown for all $k\in[0,h_{P}-1]$ and all other comparabilites are induced by transitivity.
Figure 1: The level pairs $P(k,\ell)$, $k<\ell$, which are possible in an automorphic poset $P$ of width at most three.
Given disjoint posets $P=(X,\leq_{P})$ and $Q=(Y,\leq_{Q})$, their ordinal sum is the poset $P\oplus Q:=(X\cup Y,\leq_{P}\cup\leq_{Q}\cup\;(X\times Y))$. For $P$ and $Q$ being disjoint antichains of size 2 or 3, we simply write “#P#Q\#P\#Q” as shortcut for the isomorphism type of $P\oplus Q$, e.g., $22$ indicates a 4-crown. Additionally, we write $2C$ and $3C$ for a poset consisting of two (three) disjoint chains of height 1 without any additional comparabilities between different points. The shortcuts are illustrated in Figure 1.
For a mapping $f:X\rightarrow Y$, we denote by $f|_{Z}$ the pre-restriction of $f$ to a subset $Z$ of its domain and by $f|^{Z}$ its post-restriction to a subset $Z$ of its codomain with $f[X]\subseteq Z$. For $X^{\prime}\supseteq X$, we call a mapping $f^{\prime}:X^{\prime}\rightarrow Y$ an extension of $f$, if $f^{\prime}|_{X}=f$.
An order homomorphism $r:P\rightarrow P$ is called a retraction of the poset $P$ if $r$ is idempotent, and an induced sub-poset $R$ of $P$ is called a retract of $P$ if a retraction $r:P\rightarrow P$ exists with $r[X]=R$. For the sake of simplicity, we identify $r$ with its post-restriction and write $r:P\rightarrow R$. A poset $P$ has the fixed point property iff every retract of $P$ has the fixed point property [9, Th. 4.8]. We need the following well-known result about the existence of spanning retracts:
Lemma 2.1 ([2, p. 232]).
Let $P$ be be a finite poset and $r:P\rightarrow R$ a retraction. There exists a retraction $v:P\rightarrow V$ with $V\simeq R$, $\min V\subseteq\min P$, $\max V\subseteq\max P$, and $V(1\rightarrow h_{V}-1)=R(1\rightarrow h_{R}-1)$.
A poset is called automorphic if it has a fixed point free automorphism, and it is called minimal automorphic if additionally every proper retract of it has the fixed point property. According to [9, Prop. 4.38], a poset $P$ of width at most three is automorphic iff for all level indices $0\leq k<\ell\leq h_{P}$, the level-pair $P(k,\ell)$ is one of the seven posets shown in Figure 1. The result goes back to [6, Th. 10].
3 Nice sections
3.1 Definition and basic properties
The following definition is due to Niederle [5]:
Definition 3.1.
[…]
  • $C_{j}:=c_{0,j}<\ldots<c_{h_{P},j}$ is a chain for all $j\in\{0,1,2\}$;
  • $\{c_{k,0},c_{k,1},c_{k,2}\}$ is an antichain for all $k\in[0,h_{P}]$;
  • The mapping $c_{k,j}\mapsto c_{k,(j+1)\!\!\!\mod 3}\;$ for all $k\in[0,h_{P}]$, $j\in\{0,1,2\}$, is an automorphism of $P$;
  • $P(k,k+1)\not\simeq 33$ for all level indices $k\in[0,h_{P}-1]$.
A section $P$ is called nice if $P$ does not contain an irreducible point. The symbol $\mathfrak{N}$ denotes the class of nice sections. A tower of sections is an ordinal sum of sections.
Simple examples of nice sections are the 6-crown stacks, whereas a 4-crown stack is a tower of nice sections, i.e., of 2-antichains.
For a section $P$ of width 3, the first and second property imply $c_{0,j}\lessdot\ldots\lessdot c_{h_{P},j}$ for all $j\in\{0,1,2\}$. We call the chains $C_{0},C_{1},C_{2}$ the main chains of $P$. Clearly, $P(k)=\{c_{k,0},c_{k,1},c_{k,2}\}$ for all $k\in[0,h_{P}]$. We define mappings $\lambda:P\rightarrow[0,h_{P}]$ and $\gamma:P\rightarrow\{0,1,2\}$ by setting $\lambda(c_{k,j}):=k$ and $\gamma(c_{k,j}):=j$, hence $\{p\}=P(\lambda(p))\cap C_{\gamma(p)}$ for all $p\in P$. Two consecutive levels of $P$ form a 6-crown or are of type $3C$. The level-index function $\lambda$ and the chain-index function $\gamma$ always refer to the levels and main chains of $P$; for a retract $R$ of $P$, we thus have $\{x\}=P(\lambda(x))\cap C_{\gamma(x)}$ for all $x\in R$.
Niederle [5, p. 121 and p. 125] has shown that the decomposition of a tower of sections into sections is unique, and that a poset of width at most three has not the fixed point property iff it has a proper retract being a tower of nice sections. We conclude that an automorphic poset $P$ of width at most 3 not being a tower of nice sections is not minimal automorphic. Moreover, if $P=S_{1}\oplus\cdots\oplus S_{n}$ is a tower of nice sections $S_{1},\ldots S_{n}$, it is not hard to see that $P$ is minimal automorphic iff $S_{i}$ is minimal automorphic for every $i\in[1,n]$. The objects we have to deal with in investigating minimal automorphic posets of width three are thus the nice sections of width three.
Figure 2: Two nice sections with 4-crown stacks as retracts (hollow dots). The retraction of the 6-crown stack on the left is from [3, Fig. 5.10].
In a nice section $P$ not being a 2-antichain, the level-pairs $P(0,1)$ and $P(h_{P}-1,h_{P})$ must form 6-crowns. The only nice sections of height 0, 1, and 2 are thus the 2-antichains, the 6-crowns, and the 6-crown stacks of height 2. All three are minimal automorphic. Two nice sections with height 3 are shown in Figure 2. Both contain a 4-crown stack of height 2 as retract.
3.2 Crowned sections and split levels
Our goal is a recursive approach for the determination of the nice sections of width three having a 4-crown stack as retract. However, our results will concern a larger class of sections containing the nice sections of width three as a proper subclass:
Definition 3.2.
We call a section $P$ of width three a crowned section if $P(0,1)$ and $P(h_{P}-1,h_{P})$ are 6-crowns. The symbol $\mathfrak{C}$ denotes the class of crowned sections.
For a retraction $r:P\rightarrow R$ of $P\in\mathfrak{C}$ onto a 4-crown stack, we are interested in proper lower or upper segments of $R$ which are the image of a proper lower or upper segment of $P$:
Definition 3.3.
Let $P\in\mathfrak{C}$, let $r:P\rightarrow R$ a retraction onto a 4-crown stack, and let $R(\ell)=\{a,b\}$ with $\lambda(a)\leq\lambda(b)$. We call $R(\ell)$ a down-split level if
$\ell<h_{R}\quad\>\textrm{and}\>\quad R(0\rightarrow\ell)=r[P(0\rightarrow\lambda(b))],$
and we call $R(\ell)$ an up-split level if
$0<\ell\quad\>\textrm{and}\>\quad R(\ell\rightarrow h_{R})=r[P(\lambda(a)\rightarrow h_{R})].$
In the retract of the poset on the left in Figure 2, the level set $R(0)$ is a down-split level, the level set $R(2)$ is an up-split level, and $R(1)$ is neither down- nor up-splitting. For the retract of the poset on the right of the figure, the down-split levels are $R(0)$ and $R(1)$ and the up-split levels are $R(1)$ and $R(2)$.
We agree on that for the rest of this section the poset $P$ is a crowned section and $r:P\rightarrow R$ is a retraction onto a 4-crown stack. This implies $h_{R}\geq 2$: Suppose to the contrary that $R$ is a 4-crown. For the corresponding spanning retraction $r^{\prime}:P\rightarrow R^{\prime}$, the poset $Q:=P(0,1)\cup P(h_{P})$ has the retraction $r^{\prime}|_{Q}:Q\rightarrow R^{\prime}$ onto the 4-crown $R^{\prime}$. But this is impossible because $Q$ is a 6-crown or a 6-crown stack of height 2, or it contains a point $x\in Q(1)$ with $R^{\prime}(0)<x<R^{\prime}(1)$.
In our investigation of split levels, the level sets $R(0)$, $R(1)$, $R(h_{R}-1)$, and $R(h_{R})$ frequently will play a special role. We notate the results for $R(0)$ and $R(1)$ only; the results for $R(h_{R})$ and $R(h_{R}-1)$ are dual.
We start with a simple observation: Because the down-sets $r^{-1}(a)$, $a\in R(0)$, are pairwise disjoint,
$R(0)\cap P(0)\not$$\emptyset$. (1)
Lemma 3.4.
Assume that the level set $R(\ell)$ is contained in a single level set of $P$. For $1\leq\ell\leq h_{R}-1$, the level set $R(\ell)$ is a split level, and for $\ell=0$, one of the level sets $R(0)$ and $R(1)$ is a split level.
Proof.
Let $R(\ell)$ be contained in $P(k)$ and let $z\in P(k)\setminus R(\ell)$.
  • $\ell\in[1,h_{R}-1]$: Then $k\in[1,h_{P}-1]$. Depending on $\lambda(r(z))\leq k$ or $\lambda(r(z))\geq k$, the set $R(\ell)$ is a down-split level or an up-split level.
  • $\ell=0$: Then $k=0$ due to (1). If $r(z)\in R(0)$, then $R(0)$ is a down-split level. And in the case of $r(z)\notin R(0)$, we have $R(0)<x$ or $z<x$ for all $x\in P(1\rightarrow h_{P})$ because $P$ is crowned, and $R(1)$ is an up-split level.
Lemma 2.1 and Lemma 3.4 together yield that a crowned section has a 4-crown stack as retract iff it has a 4-crown stack as retract containing a split level. On the one hand, this inconspicuous result brings our recursive approach into motion in Section 3.3. On the other hand, this approach becomes more effective and the concept of split levels more applicable if we can show that frequently there exist additional split levels not induced by the action of $r$ on $P(0)$ and $P(h_{P})$. In order to separate induced split levels from independent ones, we discriminate two cases:
  • $R(0)\subset P(0)$: Looking at the proof Lemma 3.4, we realize that if $R(0)$ is not a down-split level, then $R(1)$ is forced to be an up-split level.
  • $R(0)\not\subset P(0)$: Due to (1), there exists a point $a\in R(0)\cap P(0)$. For the point $b\in R(0)\setminus\{a\}$, we have $r[P(0)\setminus\{a\}]=\{b\}$, hence $r[P(1)\setminus\{c_{1,\gamma(b)}\}]\subseteq R(1\rightarrow h_{R})$. Nevertheless, two alternatives remain for the level set $R(1)$: it can be an up-split level or no split level at all.
In consequence, we define
Definition 3.5.
We call a split level $R(\ell)$ induced if
  • $\ell=0$ or $\ell=h_{R}$;
  • $\ell=1$, $R(0)\subset P(0)$, and $r[P(0)]\not\subset R(0)$;
  • $\ell=h_{R}-1$, $R(h_{R})\subset P(h_{P})$, and $r[P(h_{P})]\not\subset R(h_{R})$.
Otherwise, we call $R(\ell)$ independent.
For the rest of this section, we investigate structures which ensure the existence of independent split levels. We start with
Lemma 3.6.
Let $R(\ell)=\{a,b\}$ with $\lambda(a)<\lambda(b)$.
  1. 1.
    $\ell=0$: The level set $R(1)$ is an independent up-split level if
    $a$ $<P(\min\lambda[R(1)]).$
  2. 2.
    $1<\ell<h_{R}$: If
    $a$ $<P(\min\lambda[R(\ell+1)])$
    $\>\textrm{and}\>\quad b$ $>P(\max\lambda[R(\ell-1)]),$
    then $R(\ell-1)$ is a down-split level or $R(\ell+1)$ is an up-split level.
Proof.
1) We have $\lambda(a)=0$ according to (1). Due to $P(0)\setminus\{a\}<b$, we have $r(x)=b$ for all $x\in P(0)\setminus\{a\}$, thus $R(0)<r(y)$ for all $y\in P(\min\lambda[R(1)])$, and $R(1)$ is an up-split level. It is an independent one because of $R(0)\not\subset P(0)$.
2) Let $k:=\lambda(a)$, $k^{-}:=\max\lambda[R(\ell-1)]$, $j:=\lambda(b)$, and $j^{+}:=\min\lambda[R(\ell+1)]$. Let $x\in P(k)\setminus\{a\}$. If $r(x)\in\{b\}\cup R(\ell+1\rightarrow h_{R})$, then $P(k,j^{+})\simeq 33$ yields $R(\ell)<r(z)$ for all $z\in P(j^{+})$, and $R(\ell+1)$ is an up-split level.
Now assume $r(x)\in R(0\rightarrow\ell-1)\cup\{a\}$ for all $x\in P(k)\setminus\{a\}$. Then $r(z)\in R(0\rightarrow\ell-1)\cup\{a\}$ for all $z\in P(k^{-})$, and $P(k^{-},j)\simeq 33$ additionally yields $r(z)<b$, hence $r(z)\in R(0\rightarrow\ell-1)$, and $R(\ell-1)$ is a down-split level.
The retracts in Figure 2 illustrate the second part. In both, the level set $R(1)$ fulfills the assumptions, and $R(0)$ or $R(2)$ has to be a split level. In fact, both are.
Figure 3: The labeling of the points in $R(\ell,\ell+1)$ used in Lemma 3.7, Corollary 3.8, and Proposition 3.9. We assume $\lambda(a)\leq\lambda(b)$ and $\lambda(v)\leq\lambda(w)$.
For the rest of this section, we deal with a 4-crown $R(\ell,\ell+1)$. As indicated in Figure 3, we denote the points of $R(\ell)$ with $a$ and $b$ and assume $\lambda(a)\leq\lambda(b)$, and we denote the points of $R(\ell+1)$ by $v$ and $w$ and assume $\lambda(v)\leq\lambda(w)$.
Lemma 3.7.
Let $0\leq\ell\leq h_{R}-1$.
1) $P(\lambda(a),\lambda(w))\simeq 33$.
2) If $P(\lambda(b),\lambda(v))\simeq 33$, then $R(\ell)$ is a down-split level, $R(\ell+1)$ is an up-split level, and at least one of these split levels is independent.
3) If $P(\lambda(b),\lambda(v))$ is a 6-crown and $P(\lambda(b),\lambda(w))\simeq 33$, then $R(\ell)$ is a down-split level or $R(\ell+1)$ an up-split level.
Proof.
1) The points $a,c_{\lambda(a),\gamma(b)}$ and $w,c_{\lambda(w),\gamma(v)}$ form a 4-crown.
2) The assumption yields $P(\lambda(b))<R(\ell+1)$ and $R(\ell)<P(\lambda(v))$. In consequence, $R(\ell)$ is a down-split level and $R(\ell+1)$ an up-split level. For $1\leq\ell\leq h_{R}-2$, both split levels are independent, and for $\ell=0$, $R(0)$ being a down-split level makes $R(1)$ independent. For $\ell=h_{R}-1$, apply the dual argument.
3) The point $v$ has two lower covers $b$ and $z$ in $P(\lambda(b)),\lambda(v))$, and we have $z<R(\ell+1)$, hence $r(z)\in R(0\rightarrow\ell)$. For the point $y\in P(\lambda(b))\setminus\{b,z\}$, we have $y<P(\lambda((v))\setminus\{v\}$. If $r(y)\in R(0\rightarrow\ell)$, then $R(\ell)$ is a down-split level, and in the case of $r(y)\in R(\ell+1\rightarrow h_{R})$, the level $R(\ell+1)$ is an up-split level. The addendum follows directly from the definition of an independent split level.
Figure 4: Illustration of Corollary 3.8.
Corollary 3.8.
Let $1\leq\ell\leq h_{R}-2$ and let $P(\lambda(b),\lambda(v))$ form a 6-crown. Furthermore, as shown in Figure 4a, let $b^{\prime}$ be the point in $P(\lambda(b))\setminus\{b\}$ with $b^{\prime}<v$, and let correspondingly $v^{\prime}$ be the point in $P(\lambda(v))\setminus\{v\}$ with $v^{\prime}>b$.
If neither $R(\ell)$ nor $R(\ell+1)$ is splitting, then the poset induced by the points in $P(\lambda(a))\cup P(\lambda(b))\cup P(\lambda(v))\cup P(\lambda(w))$ is isomorphic to the poset shown in Figure 4b. Furthermore, $r(v^{\prime})=b$ and $r(b^{\prime})=v$.
Proof.
The relation $\lambda(a)<\lambda(b)$ is due to Lemma 3.4. Because of the dual of Lemma 3.7.3, the level pair $P(\lambda(a),\lambda(v))$ cannot be of type 33, and $P(\lambda(a),\lambda(v))$ must be a 6-crown. This implies $P(\lambda(a),\lambda(b))\simeq 3C$, and $a<v$ yields $a<b^{\prime}$. Due to Lemma 3.7.1, the type of $P(\lambda(a),\lambda(w)))$ is 33. Finally, $r(v^{\prime})\not=b$ would yield $r(v^{\prime})\in R(\ell+1\rightarrow h_{R})$, and depending on the image of the point in $P(\lambda(b))\setminus\{b,b^{\prime}\}$, the level set $R(\ell)$ would be a down-split level or $R(\ell+1)$ an up-split level. The rest is dual.
Proposition 3.9.
Let $R$ be a spanning retract and let $P(k,k+1)$ and $P(k+1,k+2)$ both be 6-crowns for a level index $k\in[1,h_{P}-3]$. With
[…]
the level set $R(\ell)$ is a down-split level or the level set $R(\ell+1)$ is an up-split level, and at least one of these split levels is independent.
Proof.
Because $R$ is spanning, the level index $\ell$ is well-defined with $\ell\leq h_{R}-1$. With our convention about the points in $R(\ell,\ell+1)=\{a,b,v,w\}$, the definition of $\ell$ implies $\lambda(a)\leq k$, $\lambda(b)\leq k+1$, $\lambda(v)\geq k+1$, and $P(\lambda(b),\lambda(v))\not\simeq 3C$.
Assume $\ell=0$. Then $\lambda(b)=0$ and $\lambda(v)\geq k+1\geq 2$ yield $P(\lambda(b),\lambda(v))\simeq 33$, and Lemma 3.7.2 delivers the result. The proof for $\ell+1=h_{R}$ is dual.
Now let $1\leq\ell\leq h_{R}-2$. In the case of $P(\lambda(b),\lambda(v))\simeq 33$, use Lemma 3.7.2 again.
Suppose thus that $P(\lambda(b),\lambda(v))$ is a 6-crown. If the type of $P(\lambda(a),\lambda(v))$ or $P(\lambda(b),\lambda(w))$ is 33, then apply Lemma 3.7.3 or its dual. Otherwise, we must have $\lambda(v)=k+1=\lambda(w)$. According to Lemma 3.4, $R(\ell+1)$ is an up-split level, hence independent, or a down-split level, which is independent for $\ell+2<h_{R}$.
Let thus $R(\ell+1)$ be a down-split level and $\ell+2=h_{R}$. Due to $R(\ell+1)\subset P(k+1)$, we have $R(h_{R})\not\subset P(k+2)$, thus $h_{P}\geq k+3$, and $R(\ell+1)<P(h_{P})$ follows. We conclude $r[P(h_{R})]=R(h_{R})$, and $R(\ell+1)$ is independent.
3.3 A recursive approach
[…]
Definition 3.10.
Let $P\in\mathfrak{C}$. We call a quadrupel $(k,D,s,t)$ a retractive down-split of $P$ if
$k$ $\in[0,h_{P}-1],$
$D$ $\subset P(k+1\rightarrow h_{P})\>\textrm{ is a down-set,}\>$
$s$ $:P(k+1\rightarrow h_{P})\setminus D\rightarrow S\>\textrm{ is a retraction,}\>$
$\>\textrm{and}\>\quad\quad\quad\quad t$ $:P(0\rightarrow k)\rightarrow T\>\textrm{ is a retraction,}\>$
$\>\textrm{where }\>S\>\textrm{ and }\>T$ are 2-antichains or 4-crown stacks,
and we call a quadrupel $(k,U,s,t)$ a retractive up-split of $P$ if
$k$ $\in[1,h_{P}],$
$U$ $\subset P(0\rightarrow k-1)\>\textrm{ is an up-set,}\>$
$s$ $:P(0\rightarrow k-1)\setminus U\rightarrow S\>\textrm{ is a retraction,}\>$
$\>\textrm{and}\>\quad\quad\quad\quad t$ $:P(k\rightarrow h_{P})\rightarrow T\>\textrm{ is a retraction,}\>$
$\>\textrm{where }\>S\>\textrm{ and }\>T$ are 2-antichains or 4-crown stacks.
In order to avoid repetitions, we use the symbols $S$ and $T$ in what follows always as in this definition: for a retractive down-split $(k,D,s,t)$ of $P\in\mathfrak{C}$, we always have $S=s[P(k+1\rightarrow h_{P})\setminus D]$ and $T=t[P(0\rightarrow k)]$, and correspondingly for a retractive up-split. In a retractive down-split $(k,D,s,t)$, the poset $T$ is thus a retract of the lower segment of $P$ ending with level $k$ and $D$ is a down-set in $P(k+1\rightarrow h_{P})$, whereas in a retractive up-split $(k,U,s,t)$, the poset $T$ is a retract of the upper segment of $P$ starting with level $k$ and $U$ is an up-set in $P(0\rightarrow k-1)$. The reader will observe that in a retractive down-split $(h_{P}-1,D,s,t)$ and a retractive up-split $(1,U,s,t)$ we must have $D\leq 1$ and $U\leq 1$, respectively, because $S$ contains at least two points.
We have seen in Section 3.2 that a crowned section has a 4-crown stack as retract iff it has such a retract with a split level. In terms of retractive splits this means
Lemma 3.11.
There exists a retraction of $P\in\mathfrak{C}_{2}$ onto a 4-crown stack iff
  • a retractive down-split $(k,D,s,t)$ of $P$ exists for which $T(h_{T})<S(0)$ and for which the mapping
    $p$ $\mapsto\begin{cases}s(p),&\>\textrm{if }\>p\in P(k+1\rightarrow h_{P})\setminus D,\\ t(p),&\>\textrm{if }\>p\in P(0\rightarrow k)\end{cases}$
    can be extended to a retraction $r:P\rightarrow T\oplus S$,
  • or a retractive up-split $(k,U,s,t)$ of $P$ exists for which $S(h_{S})<T(0)$ and for which the mapping
    $p$ $\mapsto\begin{cases}s(p),&\>\textrm{if }\>p\in P(0\rightarrow k-1)\setminus U,\\ t(p),&\>\textrm{if }\>p\in P(k\rightarrow h_{P})\end{cases}$
    can be extended to a retraction $r:P\rightarrow S\oplus T$.
We call a retraction $r$ and a retractive split matching if they are coupled in this way.
This result gives raise to a recursive approach to determine the crowned sections having a 4-crown stack as retract. Let $\mathfrak{S}$ be a subclass of $\mathfrak{C}$ closed under duality, e.g., the subclass of nice sections of width three. For a retractive split of a poset $P\in\mathfrak{S}$, we need a $t$-base $V$ and an $s$-base $W$ of a poset $P\in\mathfrak{S}$ of width 3, i.e.,
  • $V$ is a lower segment or an upper segment of $P$ having a retraction $t$ onto a 2-antichain or a 4-crown stack $T$.
  • $W$ is an upper segment of $P$ having a retraction of $W\setminus D$ onto a 2-antichain or a 4-crown stack $S$ for a down-set $D\subset W$, or $W$ is a lower segment of $P$ having a retraction of $W\setminus U$ onto a 2-antichain or a 4-crown stack $S$ for an up-set $U\subset W$.
  • $V$ and $W$ are coupled by $V\cap W=\emptyset$ and $P=V\cup W$.
Assume that we have listed all lower segments up to height $n$ of sections contained in $\mathfrak{S}$ which have a 2-antichain or a 4-crown stack as retract. Given a poset $P\in\mathfrak{S}$ of height $n+1$, we can quickly identify all candidates for $t$-bases in $P$ by checking which of the lower segments in our list are lower segments $P(0\rightarrow k)$ of $P$ and which of their duals are upper segments $P(k\rightarrow n+1)$ of $P$. Of course, $P(0)$ and $P(n+1)$ are always candidates. The poset $P$ has a 4-crown stack as retract iff one of the corresponding segments $P(k+1\rightarrow n+1)$ or $P(0\rightarrow k-1)$ provides an $s$-base which can be combined with the $t$-base as described in Lemma 3.11.
In order to test the approach we need a sub-class of $\mathfrak{C}$ combining plenty of interesting posets with good structural access. We choose the class of the nice sections with horizon two. We introduce them in Section 4, and in Section 5 we identify all of them with height up to six having a 4-crown stack as retract.
4 Nice sections with horizon two
4.1 Definition and isomorphism types
For a crowned section $P\in\mathfrak{C}$ with height $h_{P}\geq 2$, we have $P(0,h_{P})\simeq 33$. There exists thus a smallest integer $\eta$ with
$P(k,k+\eta)\simeq 33\quad\>\textrm{for all }\>k\in[0,h_{P}-\eta].$
We call this integer the horizon of $P$. It is the distance in which all details become blurred. The sections in Figure 2 have horizon two.
For an integer $\eta\in\mathbb{N}$, the symbols $\mathfrak{C}_{\eta}$ and $\mathfrak{N}_{\eta}$ denote the classes of crowned sections and nice sections with horizon $\eta$ and height at least 2. The set $\mathfrak{C}_{1}$ is empty, and the only crowned sections not contained in $\mathfrak{C}_{\eta}$ for any $\eta\geq 2$ are the 6-crowns. For the rest of this article we focus on $\eta=2$, and for this horizon, we can skip the discrimination between nice sections and crowned sections:
Proposition 4.1.
$\mathfrak{C}_{2}=\mathfrak{N}_{2}$.
Proof.
Let $P\in\mathfrak{C}_{2}$. We have to prove that $P$ does not contain an irreducible point. Let $x\in P(k)$, $k\in[0,h_{P}-1]$. For the upper covers of $x$, only the case $P(k,k+1)\simeq 3C$ is of interest which yields $k\leq h_{P}-2$. One upper cover of $x$ is contained in $P(k+1)$, and depending on the type of $P(k+1,k+2)$, the layer $P(k+2)>x$ contains one or even two additional upper covers of $x$. The argument for lower covers is dual.
The 6-crown stacks of height greater than one belong to $\mathfrak{N}_{2}$, and Farley [3, Prop. 4.1] showed that all 6-crown stacks with equal height are isomorphic. The following theorem generalizes this result:
Theorem 4.2.
The isomorphism type of a poset $P\in\mathfrak{N}_{2}$ is uniquely determined by the types of its level-pairs $P(k,k+1)$, $k\in[1,h_{P}-2]$. In particular, there exist $2^{n-2}$ isomorphism types of posets of height $n\geq 2$ in $\mathfrak{N}_{2}$. Furthermore, every permutation of $P(0)$ can be uniquely extended to an automorphism of $P$.
Proof.
For each integer $n\geq 2$ and each index set $K\subseteq\{0,\ldots,n-1\}$ with $0,n-1\in K$, we can construct a poset $P\in\mathfrak{N}_{2}$ with $h_{P}=n$ and $P(k,k+1)$ being a 6-crown iff $k\in K$: We start with the union $Q$ of three pairwise disjoint chains of length $n$, extend $Q(k,k+1)$ to an arbitrary 6-crown for every $k\in K$, and complete the partial order relation by adding $Q(k)\times Q(\ell)$ for all $k,\ell\in[0,n]$ with $k\leq\ell-2$. The resulting section is an element of $\mathfrak{C}_{2}$ and belongs thus to $\mathfrak{N}_{2}$ (Proposition 4.1).
Now let $P=(X,\leq_{P})$ and $Q=(Y,\leq_{Q})$ be elements of $\mathfrak{N}_{2}$ and let $\alpha:P(0)\rightarrow Q(0)$ be a bijection. We define on $X$ and $Y$ the relations
[…]
where $\upsilon$ is the level-function of $Q$. It is easily seen that we can extend $\alpha$ to an isomorphism $\beta:(X,\prec_{P})\rightarrow(Y,\prec_{Q})$ iff $h_{P}=h_{Q}$ and
$\forall\;k\in[1,h_{P}-2]:\quad P(k,k+1)\>\textrm{ 6-crown}\>\;\;\Leftrightarrow\;\;Q(k,k+1)\>\textrm{ 6-crown,}\>$
and that in this case $\beta$ is uniquely determined by $\alpha$. Every isomorphism from $(X,\prec_{P})$ to $(Y,\prec_{Q})$ sends $P(k)$ onto $Q(k)$ for all $k\in[0,h_{P}]$, and because $P$ and $Q$ have horizon 2, the isomorphisms between $(X,\prec_{P})$ and $(Y,\prec_{Q})$ are exactly the isomorphisms between $P$ and $Q$.
Due to this result, we can always draw the Hasse-diagram of a poset $P\in\mathfrak{N}_{2}$ in a standardized way: all level-pairs $P(k,k+1)$ being 6-crowns are drawn with a “void” in the center as in Figure 2.
We denote the classes of segments, lower segments, and upper segments of nice sections with horizon 2 by $\mathfrak{N}^{Seg}_{2},\mathfrak{N}^{LS}_{2}$, and $\mathfrak{N}^{US}_{2}$, respectively. Theorem 4.2 extends to these classes: Two segments of equal height $n$ are isomorphic iff both are 3-antichains (the case $n=0$) or all of their consecutive level-pairs are isomorphic (the case $n\geq 1$), and also the extension of a level permutation to an automorphism is always possible.
4.2 Properties of retracts
We start our investigation with three simple observations:
Corollary 4.3.
Let $P\in\mathfrak{N}^{Seg}_{2}$ and let $r:P\rightarrow R$ be a retraction.
1) For all $\ell\in[0,h_{R}]$ there exists an index $k\in[0,h_{P}-1]$ with $R(\ell)\subset P(k,k+1)$. In particular, $R(0)\subset P(0,1)$ if $R(0)\geq 2$.
2) Let $\ell\in[0,h_{R}]$ with $R(\ell)=2$ and $0<\rho:=\max\lambda[R(\ell)]<h_{P}$. We have $r[P(\rho+1\rightarrow h_{P})]=R(\ell+1\rightarrow h_{R})$ if
$r^{-1}(a)\cap P(\rho-1)\not=\emptyset\quad\>\textrm{for all }\>a\in R(\ell).$ (2)
In particular, the level set $R(\ell+1)$ is an up-split level.
Proof.
The first proposition follows from the definition of the horizon and (1). For the second one, observe that the assumption implies $a\leq r(y)$ for all $a\in R(\ell)$ and all $y\in P(\rho+1)$, hence $R(\ell)<r(y)$.
Figure 5: The construction in the proof of Corollary 4.4.
Corollary 4.4.
Let $P\in\mathfrak{N}^{Seg}_{2}$ with $h_{P}\geq 2$ and $P(0,1)\simeq 3C$. P has a retract $R$ with $R(0,1)$ being a 4-crown iff $P(2\rightarrow h_{P})$ has a retract $W$ with $W(0)$ being a 2-antichain.
Proof.
$\Rightarrow$”: Let $R(0)=\{a,b\}$ with $\lambda(a)\leq\lambda(b)$. Due to the first part of Corollary 4.3, we have $R(0)\subset P(0,1)$. In the case of $\lambda(b)=1$, the equation $r[P(2\rightarrow h_{P})]=R(1\rightarrow h_{P})$ follows with (2). In the case of $\lambda(b)=0$, the relation $R(0)<P(2\rightarrow h_{P})$ enforces $r[P(2\rightarrow h_{P})]\subseteq R(1\rightarrow h_{P})$, and $P(0,1)\simeq 3C$ yields $R(1)\subset P(2\rightarrow h_{P})$.
$\Leftarrow$”: If $w:P(2\rightarrow h_{P})\rightarrow W$ is a retraction with $W(0)$ being a 2-antichain, select two points $a,b\in P(0)$ and define a retraction $r:P\rightarrow W\cup\{a,b\}$ by (cf. Figure 5)
$r(x)$ $:=\begin{cases}w(x),&\>\textrm{if }\>x\in P(2\rightarrow h_{P}),\\ a,&\>\textrm{if }\>x\in\{a,c_{1,\gamma(a)}\},\\ b,&\>\textrm{if }\>x\in P(0,1)\setminus\{a,c_{1,\gamma(a)}\}.\end{cases}$
4.3 Retractive up- and down-splits
Let $P\in\mathfrak{N}_{2}$ and let $r:P\rightarrow R$ be a retraction onto a 4-crown stack $R$. Due to the horizon of $P$, every level set of $R$ is either contained in a single level set of $P$ or consists of points belonging to consecutive level sets of $P$. Now Lemma 3.4 and Lemma 3.6 tell us that every level-set of $R$ gives raise to matching retractive split of $P$. On the other hand, for a given retractive split of $P$ it is easy to check if there exists a matching retraction of $P$ onto a 4-crown stack:
Theorem 4.5.
Let $P\in\mathfrak{N}_{2}$. For a retractive down-split $(k,D,s,t)$ of $P$, the existence of a matching retraction $r:P\rightarrow T\oplus S$ is equivalent to
$\begin{split}T(h_{T})&<_{P}S(0),\\ \forall\;d\in D\;\exists\;v\in T(h_{T}):\;p&\not<_{P}d\>\textrm{ for all }\>p\in t^{-1}(v).\end{split}$ (3)
which implies $D\subseteq P(k+1)$, and for a retractive up-split $(k,U,s,t)$ of $P$, there exists a matching retraction $r:P\rightarrow S\oplus T$ iff
$\begin{split}S(h_{S})&<_{P}T(0),\\ \forall\;u\in U\;\exists\;v\in T(0):\;u&\not<_{P}p\>\textrm{ for all }\>p\in t^{-1}(v),\end{split}$ (4)
which implies $U\subseteq P(k-1)$.
Proof.
Let $r:P\rightarrow R$ be a retraction onto a 4-crown stack $R$ matching a retractive up-split $(k,U,s,t)$. Due to $r[P(k\rightarrow h_{P})]=T$, the second part of Condition (4) is necessary for $r[U]\subseteq T$. The addendum follows because of $P(k-2)<_{P}P(k)\supset T(0)$.
Now assume that (4) holds for $(k,U,s,t)$ and let $u\in U\subseteq P(k-1)$. There exists a point $\tau(u)\in T(0)$ with $u\not<_{P}p$ for all $p\in t^{-1}(\tau(u))$. We define $\rho(u)$ as the single point contained in $T(0)\setminus\{\tau(u)\}$. We define a mapping $r:P\rightarrow P$ by setting for $x\in P$
$r(x)$ $:=\begin{cases}s(x),&\>\textrm{if }\>x\in P(0\rightarrow k-1)\setminus U,\\ \rho(x),&\>\textrm{if }\>x\in U,\\ t(x),&\>\textrm{if }\>x\in P(k\rightarrow h_{P}),\end{cases}$
The mapping $r$ is clearly idempotent. In order to show that it is order-preserving, let $x<_{P}y$. Only three cases have to be discussed:
  • $x\in P(0\rightarrow k-1)\setminus U$ and $y\in U\cup P(k\rightarrow h_{P})$ yields $r(x)\in S<_{S\oplus T}T\ni r(y)$.
  • $x\in U$, $y\in P(0\rightarrow k-1)$ is not possible due to $U\subseteq P(k-1)$.
  • $x\in U$ and $y\in P(k\rightarrow h_{P})$: $x\not<p$ for all $p\in t^{-1}(\tau(x))$ delivers $t(y)\in T\setminus\{\tau(x)\}$. Because $\rho(x)$ is the only minimal point of this poset, $r(x)=\rho(x)\leq t(y)=r(y)$.
Figure 6: Illustrations for the proof of Corollary 4.6. The three levels show $P(k-1\rightarrow k+1)$ for choices $b\in P(k-1)$, $w\in P(k+1)$ (which are not mandatory). The point $z$ is not mapped to $a$ by the retraction $s$ of $P(0\rightarrow k-1)$.
Corollary 4.6.
Let $k\in[1,h_{P}-1]$ and assume that we have retractions $s:P(0\rightarrow k-1)\rightarrow S$ and $t:P(k+1\rightarrow h_{P})\rightarrow T$ with $S$ and $T$ being 2-antichains or 4-crown stacks with
$s^{-1}(a)$ $=\{a\}\>\textrm{ for a point }\>a\in S(h_{S}),$
$\>\textrm{and}\>\quad t^{-1}(v)$ $=\{v\}\>\textrm{ for a point }\>v\in T(0).$
The 4-crown stack $S\oplus T$ is a retract of $P$ if $P(k-1\rightarrow k+1)$ is not a 6-crown stack.
Proof.
Let $S(h_{S})=\{a,b\}$ and $T(0)=\{v,w\}$. The assumptions imply $a\in P(k-1)$ and $v\in P(k+1)$. Assume $P(k,k+1)\simeq 3C$ (the other case is dual). According to Lemma 2.1, we can assume $S(h_{S})\subset P(k-1)$ without loss of generality. Let $P(k-1)=\{a,b,z\}$.
Let $U\subset P(k)$ be the set of upper covers of $a$ in $P(k)$. Applying Theorem 4.2 on $P(0\rightarrow k)$, we see that without loss of generality we can assume that none of the points in $U$ is below $v$ (Figure 6).
No point of $P(k)\setminus U$ is above $a$ and the point $b$ is the only maximal point of $S\setminus\{a\}$. Due to $s^{-1}(a)=\{a\}$, the mapping $s^{\prime}:P(0\rightarrow k)\setminus U\rightarrow S$ defined by
$s^{\prime}(x)$ $:=\begin{cases}s(x),&\>\textrm{if }\>x\in P(0\rightarrow k-1),\\ b,&\>\textrm{if }\>x\in P(k)\setminus U\end{cases}$
is a retraction, and due to $t^{-1}(v)=\{v\}$, the quadrupel $(k+1,U,s^{\prime},t)$ is a retractive up-split fulfilling (4).
5 Application
With the recursive approach described in Section 3.3, we will now identify all posets $P\in\mathfrak{N}^{LS}_{2}$ with height up to six which have a 4-crown stack as retract. For some of them, e.g. the 6-crown stacks, we already know the result. We include them in the investigation because our purpose is to test our approach with as many nice sections as possible.
We encode the lower segments not being 3-antichains by the sequence of their level-pair types (cf. Theorem 4.2). The letter “1” indicates a 6-crown, “0” a type $3C$. The lower segment 111 is thus the 6-crown stack of height 3, whereas 110 indicates the lower segment starting with two 6-crowns followed by a single $3C$.
1 n
11 n 10 y
111 y 101 y 110 n 100 n
$11$ 0,3 n 1101 0 n 1011 0,2,3 n 1001 0,2 y
1110 n 1100 n $10$ y 1000 y
11111 0,3 n 11101 0,3 y 11011 0 n 10111 0,2,3 y
11001 0 y 10101 0,2,3,4 y 10011 0,2,4 y 10001 0,2,4 y
11110 y 11100 y 11010 n 10110 y
11000 n 10100 y 10010 n 10000 n
$111$ 0,3 y 111101 0,3,5 y 111011 0,3,5 n 110111 0 n
101111 0,2,3,5 y 111001 0,3,5 n 110101 0 n $101$ 0,2,3,5 y
110011 0,5 n 101011 0,2,3,4,5 n 100111 0,2,4,5 n 110001 0 n
101001 0,2,3,4,5 y 100101 0,2,4 y 100011 0,2,4,5 n 100001 0,2,4 y
111110 n 111100 n 111010 n $110$ n
101110 n 111000 n 110100 n 101100 n
110010 n 101010 y 100110 y 110000 n
101000 y $100$ y 100010 y 100000 y
Table 1: The lower segments in $\mathfrak{N}^{LS}_{2}$ with height from one to six. Explanation in text.
The lower segments with height from one to six are listed in Table 1. The letters “y” and “n” indicate the result of our investigation whether the segment has a 4-crown stack as retract or not. For the posets 1, 10, 11, 111, and 101, we refer to common knowledge. The lower segment $10$ even has a retraction $r$ to a 4-crown with $r^{-1}(v)=\{v\}$ for a point $v$ from the top level of the 4-crown, as shown in Figure 7a.
For the posets with a final 0, we can apply the dual of Corollary 4.4 on previous results in Table 1. As an example, the posets $110$ and $100$ do not have a 4-crown stack as retract, because the 6-crown $P(0,1)$ does not have a suitable retract. And $1010$ and $1000$ have a 4-crown stack as retract because 10 is marked with “y”.
The bulk of the work concerns thus the lower segments $P$ with a 6-crown as final level-pair. They all are elements of $\mathfrak{N}_{2}$. For them, the integers in Table 1 indicate the levels $k\in[0,h_{P}-1]$ for which $P(0\rightarrow k)$ can be a $t$-base according to the results for lower segments with less height. For $1111$ it is 0 and 3: 0, because the antichain $P(0)$ is always a candidate for a $t$-base, and 3, because the poset $1111(0\rightarrow 3)=111$ is marked with “y” in the previous row. Because 1 and 11 are marked with “n”, the level indices 1 and 2 are missing in the list for $1111$.
Figure 7: a) The lower segment $10$ and a retraction $r$ onto a 4-crown with $r^{-1}(v)=\{v\}$ for a point $v$ from the top level of the 4-crown. b) Illustrations for the case discrimination in the proof of Criterion 5.1.2. The hollow dots belong to $S(0)$.
The following criteria will be useful in our investigation:
Criteria 5.1.
Let $P\in\mathfrak{N}_{2}$ with $h_{P}\geq 3$. Assume that there exists a matching retraction $r$ for the retractive down-split $(k,D,s,t)$.
  1. 1.
    $k=0$ implies that $D\subset P(1)$ contains at most a single point.
  2. 2.
    $k=0$ and $P(1,2)$ being a 6-crown implies that the poset $P(3\rightarrow h_{P})$ has a 2-antichain or a 4-crown stack as retract.
  3. 3.
    $k=h_{P}-2$ implies that $D\subseteq P(k-1)$ contains at least two points.
  4. 4.
    $k=h_{P}-2$ and $P(h_{P}-2,h_{P}-1)$ being a 6-crown implies that the poset $P(0\rightarrow h_{P}-3)$ has a 2-antichain or a 4-crown stack as retract.
  5. 5.
    $k=h_{P}-1$ implies that the poset $P(0\rightarrow h_{P}-3)$ has a 2-antichain or a 4-crown stack as retract.
Proof.
1. Trivial.
2. We discriminate three cases (cf. Figure 7b). If $S(0)$ contains a point $a\in P(1)$, then it has to contain a point $b\in P(2)$, too, because of the first condition in (3). At least one of the points in $P(1)\setminus\{a\}$ has to be mapped to $b$ (first criterion), thus $s[P(3\rightarrow h_{P})]=S(1\rightarrow h_{S})$ due to (2).
Now assume that $S(0)$ does not contain a point of $P(1)$. If $S(0)$ contains a point $v\in P(3\rightarrow h_{P})$, then at least two points in $P(1)$ are below $S(0)$ in contradiction to the first criterion. And in the case of $S(0)\subset P(2)$, exactly one point $x\in P(1)$ is below $S(0)$ and has to be sent to $P(0)$. From the two remaining points in $P(1)$ each is under a single different point of $S(0)$. Due to the first criterion, they have to be mapped onto $S(0)$, and $s[P(3\rightarrow h_{P})]=S(1\rightarrow h_{S})$ follows again with (2).
3. The poset $P(h_{P}-1,h_{P})\setminus D$ has to be disconnected.
4. According to the third criterion, the set $D\subset P(h_{P}-1)$ contains at least two points. The second condition in (3) says that every point in $D$ is comparable with a single point of $T(h_{T})$ only which implies $T(h_{T})\subset P(h_{P}-2)$. A matching retraction maps $D$ onto $T(h_{T})$, and the dual of (2) yields the result.
5. The first condition in (3) yields $T(h_{T})\not\subset P(h_{P}-1)$. Now the dual of (2) yields $t[P(0\rightarrow h_{P}-3)]=T(0\rightarrow h_{T}-1)$.
The rightmost poset in Figure 9 confirms that we can neither drop the condition “$P(1,2)$ is a 6-crown” in Criterion 5.1.2 nor the condition $P(h_{P}-2,h_{P}-1)$ is a 6-crown” in Criterion 5.1.4.
Figure 8: The first three diagrams are illustrations for the proof of Lemma 5.2. Additionally the rightmost diagram shows for $P=001$ a retraction of $P\setminus D$ to a 4-crown with a singleton $D\subset P(0)$.
Lemma 5.2.
Let $P=011$. If $r:P\setminus D\rightarrow R$ is a retraction onto a 4-crown stack or a 2-antichain for a down-set $D$, then $D$ contains at least two points of $P(0)$.
Proof.
Assume that a retraction exists for $D=\{d\}$ with $d\in P(0)$. Figure 8a shows $P\setminus\{d\}$. Because $r$ cannot be extended to $P$, both upper covers $x$ and $y$ of $d$ have to be mapped to different points of $R(0)$. The encircled set ${\downarrow\,}x$ has thus to be mapped to a single point in $R(0)$. The resulting poset is shown in Figure 8b. Clearly, it does not have a 4-crown stack or a 2-antichain as retract.
Now assume $D=\{d,y\}$ with $d\in P(0)$ and $d\lessdot y$. The poset $P\setminus D$ is shown in Figure 8c. If a retraction as described in the lemma exists, it cannot be extended to $P\setminus\{d\}$. The upper covers $a$ and $b$ of $y$ are thus mapped to different points of $R(0)$. But this is impossible because of $P(0)\setminus\{d\}<\{a,b\}$.
Figure 9: The posets $1111$, $1101$, $1011$, and $1001$. As already mentioned at the end of the introduction, the poset $1011$ is also treated in [9, p. 99] and mentioned in [3, p. 136] as a very complicated case. For the poset $1001$, a retraction onto a 4-crown tower is indicated resulting from a retractive down-split $(2,D,s,t)$ based on the retraction $t$ of $P(0\rightarrow 2)=10$ shown in Figure 7a.
Now we can start with the posets of height four. They are shown in Figure 9.
  • P = 1111: Criterion 5.1.5 prevents $P(4)$ from being a candidate for an $s$-base because the 6-crown $P(0,1)$ does not have a suitable retract. Furthermore, selecting $P(1\rightarrow 4)$ as $s$-base does not work due to Criterion 5.1.2, because $P(3,4)\simeq 1$ is marked with “n” in Table 1. $1111$ does thus have no retractive down-split fulfilling (3), and because the poset is self-dual, we conclude that it does not have a 4-crown stack as retract.
  • $P=1101$: Again, $P(1\rightarrow 4)$ fails as an $s$-base due to Criterion 5.1.2. Therefore, $1101$ does not have a retractive down-split fulfilling (3).
  • $P=1011$: $P(1\rightarrow 4)$, $P(3,4)$, or $P(4)$ being a successful $s$-base is prevented by Criterion 5.1.1 together with Lemma 5.2, Criterion 5.1.4, and Criterion 5.1.5, respectively. Also $1011$ does thus not have a retractive down-split fulfilling (3). Therefore, neither $1101$ nor $1011$ has a 4-crown stack as retract.
  • $P=1001$: A retraction to a 4-crown stack resulting from a retractive down-split $(2,D,s,t)$ is indicated in Figure 9 on the right.
We continue with the posets of height five:
  • $P=11111$: Due to the Criteria 5.1.2 and 5.1.4, the upper segments $P(1\rightarrow 5)$ and $P(4,5)$ cannot be work as $s$-bases. The self-dual poset has thus no 4-crown stack as retract.
  • $P=11101$: see $P=10111$.
  • $P=11011$: Criterion 5.1.2 prevents $P(1\rightarrow 5)$ from being a successful $s$-base. Because $11011$ is self-dual, the poset cannot have a 4-crown stack as retract.
  • $P=10111$: Figure 2 shows that the 6-crown stack $Q:P(2\rightarrow 5)=111$ has a retraction $t:Q\rightarrow T$ with $T$ being a 4-crown stack and $t^{-1}(a)=\{a\}$ for a point $a\in T(0)$. With $s:P(0)\rightarrow P(0)$ being a mapping such that $s[P(0)]$ is a 2-antichain, apply Corollary 4.6 with $k=1$.
  • $P=11001$: see $10011$.
  • $P=10101,10001$: Let $s:P(0\rightarrow 2)\rightarrow S$ and $t:P(3\rightarrow 5)\rightarrow T$ retractions onto 4-crowns as shown in Figure 7a and its dual. Without loss of generality (Theorem 4.2) we assume that the single point in $P(2)\cap S$ is below the single point in $P(3)\cap T$. Now $(3,\emptyset,s,t)$ is a retractive down-split fulfilling (3).
  • $P=10011$: Let $Q:=1001$. The retraction $t$ of $Q$ indicated in Figure 9 on the right and a self-mapping $s$ of $P(5)$ onto a 2-antichain can be combined to a retractive down-split $(4,\emptyset,s,t)$ of $P$ fulfilling (3).
For some of the posets of height six, a quick positive decision is possible:
  • $P=111111,111101,101111,101101$: Combine the retractions in Figure 2.
  • $P=101001,100001$: Let $P$ be one of these posets. With $Q:=P(3\rightarrow 6)=001$, select a point $d\in Q(0)$ and let $s$ be the retraction of $Q\setminus\{d\}$ onto a 4-crown $S$ shown in Figure 8d with $a\in S\cap Q(0)$ as in the figure. Furthermore, let $t$ be the retraction of $P(0\rightarrow 2)=10$ onto a 4-crown $T$ as shown in Figure 7a with $v\in T\cap P(2)$ as in the figure. Applying Theorem 4.2 on $P(0\rightarrow 2)$, we can assume $v\lessdot a$ and $v\not<d$ and get a retractive down-split $(2,\{d\},s,t)$ fulfilling (3).
  • $P=100101$: dual to $101001$.
For $P=110011$, the Criteria 5.1.2 and 5.1.5 prevent $P(1\rightarrow 6)$ and $P(6)$ from being successful $s$-bases; the self-dual poset has thus now 4-crown stack as retract. The remaining posets of height six are $111011$, $111001$, $101011$, $110001$, and their duals. They do not have a 4-crown stack as retract:
Lemma 5.3.
None of the posets $111011$, $111001$, $101011$, and $110001$ has a 4-crown stack as retract.
Proof.
Firstly, let $P\in\mathfrak{N}_{2}$ be any nice section of height six and $R$ a retract of $P$ being a 4-crown stack. Every level set of $R$ has to be contained in a single level set or in two consecutive level sets of $P$, and we conclude that at least one of the posets $P(0\rightarrow 3)\cap R$ and $P(3\rightarrow 6)\cap R$ is a 4-crown stack.
Now let $P$ be any of the four posets and $r$ a retraction onto a 4-crown stack $R$. With $S:=P(3\rightarrow 6)\cap R$, we show in a first step that $S$ cannot be a 4-crown stack.
$P(0\rightarrow 3)=111,110$: $P(3\rightarrow 6)$ does not have a 4-crown stack as retract. $S$ being a 4-crown stack requires thus a point $x\in P(3)$ with $v:=r(x)\in P(0\rightarrow 2)=11$. Let $\{v,w\}$ be the level set of $R$ containing $v$. $x\in P(3)$ yields $r[P(0,1)]\leq v$, hence $w\notin P(0,1)$. But then, $r(z)<\{v,w\}$ holds for at least five of the points $z\in P(0,1)$, and $P(0,1)$ cannot be mapped onto a 2-antichain.
$P=101011$: $S$ being a 4-crown stack requires due to Lemma 5.2 two points $x,y\in P(3)$ to be mapped into $P(0\rightarrow 2)=10$. The points $r(x)$ and $r(y)$ belong thus to $R(0,1)$. Three cases are possible, each leads to a contradiction:
  • $\{r(x),r(y)\}$ is a level set of $R$: Let $z\in P(2)$ be the common lower cover of $x$ and $y$. Due to $P(0)<z$, the point $r(z)$ is the only minimal point of $R$.
  • $r(x)<r(y)$: Then $r(x)\in R(0)$, and $P(0)<x$ yields that $r(x)$ is the only minimal point of $R$.
  • $v:=r(x)=r(y)$: Let $\{v,w\}$ be the level set of $R$ containing $v$. Due to $r[P(0\rightarrow 2)]\leq v$, the point $w$ must belong to $P(3\rightarrow h_{P})$, and due to $\lambda(v)\leq 2$, we must have $w\in P(3)$ and $v\in P(2)$. But then $w\in S\not\ni v$.
$S$ is thus not a 4-crown stack. This requires a level set $R(\ell)=\{a,b\}$ with $a\in P(2)$ and $b\in P(3)$. Clearly, $\ell=1$. Due to $P(1)<b$, no point of $P(1)$ can be mapped to $a$. And if a point of $P(1)$ is mapped to $b$, (2) yields $R(2\rightarrow h_{R})$ being a retract of $P(4\rightarrow 6)=11$ which is impossible. But if no point of $P(1)$ can be mapped to $R(1)$, then $r[P(0,1)]=R(0)$ which is impossible, too.